Hello out there,I had some problems that I need help with,and thanks to whoever can help,The first one is,Write an equation in slope-intercept form of a line with the following characteristics:perpendicular to the graph of:2x-3y=12 and passes through (4,3).Second one is,Determine the type of relationship a scatter plot of the data would show:the distance from an object versus its apparent size.(I really don't understand that one or the next)The next one is,Determine the type of relationship plot of the data would show:the number of students attending a basketball game versus the noise level in the arena.And the last one is,Determine the x-intercept of the graph:2x+4y=8.Thanks to whoever can help me.
Answer by stanbon(48558)

Write an equation in slope-intercept form of a line with the following characteristics:perpendicular to the graph of:2x-3y=12 and passes through (4,3).
--------
The given equation has slope = 2/3
The equation you want must have slope--3/2 and pass thru (4.3)
3=(-3/2)4+b
b=9
EQUATION: y=(-3/2)x+9
-----------------
Second one is,Determine the type of relationship a scatter plot of the data would show:the distance from an object versus its apparent size.
--------
As the distance from an object gets greater the apparent size gets smaller.
So the scatter plot would have a linear appearance with a negative slope.
================
Determine the type of relationship plot of the data would show:the number of students attending a basketball game versus the noise level in the arena.
As the number of students attending increases the noise level increases.
So the scatter plot would have a linear appearance with a positive slope.
==================
And the last one is,Determine the x-intercept of the graph:2x+4y=8.
---
Let y=0, then x=4
That is the x-intercept.
==============
Cheers,
Stan H.